OpenStudy (anonymous):

a fire burned a square-shaped forest 45 miles on a side, First,draw a sketch and label the lengths of the sides then find the area.

OpenStudy (anonymous):

A = a^2 = 45^2 = 2025 mi^2

OpenStudy (anonymous):

thankss